# NeuRes: Learning Proofs of Propositional Satisfiability

### v1

@@ 1 Introduction
Boolean satisfiability (SAT) is a fundamental problem in computer science. For theory, this stems from SAT being the first problem proven NP-complete (Cook, 1971). For practice, this is due to many highly-optimized SAT solvers being used as flexible reasoning engines in a variety of tasks such as model checking (Clarke et al., 2001; Vizel et al., 2015), software verification (de Moura & Bjørner, 2011), planning (Kautz & Selman, 1992), and mathematical proof search (Heule & Kullmann, 2017). Recently, SAT has also served as a litmus test for assessing the symbolic reasoning capabilities of neural models and a promising domain for neuro-symbolic systems (Selsam et al., 2019; Selsam & Bjørner, 2019; Amizadeh et al., 2019; Cameron et al., 2020; Ozolins et al., 2022). So far, neural models only provide limited justifications (certificates/witnesses) for their unsatisfiability predictions, mostly in form of an unsatisfiable core of the formula. Verifying these certificates can be as hard as solving the original problem traditionally, which severely limits the usefulness of neural methods in a domain where correctness is critical. Therefore, we propose a neuro-symbolic model that utilizes resolution at its core to solve SAT problems by generating easy-to-check certificates.
@@
A resolution proof is a sequence of case distinctions, each involving two clauses, that ends in the empty clause (falsum). This technique can also be used to prove satisfiability by exhaustively applying it until no further new resolution steps are possible and the empty clause has not been derived. Generating such a proof is an interesting problem from a neuro-symbolic perspective because unlike other discrete combinatorial problems that have been considered before (Vinyals et al., 2015; Bello et al., 2017; Khalil et al., 2017; Kool et al., 2019; Cappart et al., 2023), it requires selecting compatible pairs of clauses from the dynamically growing input formula, as newly derived clauses are naturally considered for derivation in subsequent steps. To handle unsat (unsatisfiable) formulas, we devise three attention-based mechanisms to perform this pair-selection needed for generating resolution proofs. To handle sat (satisfiable) formulas, we augment the architecture with an assignment decoding perceptron that assigns a truth value to each literal node. We hypothesize that, despite their final goals being in complete opposition, resolution and sat assignment finding can form a mutually beneficial collaboration. On the one hand, clauses derived by resolution incrementally inject additional information into the network which allows the assignment decoder to re-calibrate its predictions with each new resolution inference. For example, deriving a single-literal clause by resolution directly implies that literal should be true in any possible sat assignment. On the other hand, finding a sat assignment absolves the resolution network from having to prove satisfiability by exhaustion. On that basis, given an input formula, NeuRes proceeds in two parallel tracks: (1) finding a sat assignment, and (2) deriving a resolution proof of unsatisfiability. At each step, our model produces both a new clause by resolution and a variable assignment. Both tracks operate on a shared representation of the problem state. Depending on which track succeeds, NeuRes produces the corresponding SAT verdict which is guaranteed to be sound by virtue of its certificate-based design. Since both of our certificate types are easy-to-check, we can afford to perform these checks at each step.
Our contributions.
@@
    We introduce three novel architectures for generating resolution proofs for CNF formulas (Section 4) and demonstrate their ability to generalize from a small set of teacher proofs (Section 6.1).
@@
    We study our model’s ability to shorten teacher proofs and devise a bootstrapped training procedure where our model progressively supplants the teacher proofs with much shorter versions (Section 6.3). This optimizes the storage consumption of resolution proofs and boosts the model’s overall performance.
@@
    We devise a resolution-aided SAT solver trained on teacher assignments and proofs, and show that it outperforms the NeuroSAT baseline by a wide margin while requiring far less training samples (Section 7).
@@ 2 Related Work
SAT Solvers and Certificates.
We refer to the annual SAT competitions (Balyo et al., 2023) for a comprehensive overview on the ever-evolving landscape of SAT solvers, benchmarks, and proof checkers. SAT solvers are complex systems with a documented history of bugs (Brummayer et al., 2010; Järvisalo et al., 2012), hence proof certificates have been partially required in this competition since 2013 (Balint et al., 2013). Unlike satisfiable formulas, there are several ways to certify unsatisfiable formulas (Heule, 2021). Resolution proofs (Zhang & Malik, 2003; Goldberg & Novikov, 2003) are easy to verify (Darbari et al., 2010), but non-trivial to generate from modern solvers based on the paradigm of conflict-driven clause learning (Marques-Silva et al., 2021). Clausal proofs, e.g., in DRAT format (Wetzler et al., 2014), are easier to generate and space-efficient, but hard to validate. Verifying the proofs can take longer than their discovery (Heule et al., 2014) and requires highly optimized algorithms (Lammich, 2020).
@@
NeuroSAT (Selsam et al., 2019) was the first study of the Boolean satisfiability problem as an end-to-end learning task. Building upon the NeuroSAT architecture, (Selsam & Bjørner, 2019) trained a simplified version to predict unsatisfiable cores and successfully integrated it in a state-of-the-art SAT solver. (Cameron et al., 2020) showed that both the NeuroSAT architecture and a newly introduced deep exchangeable architecture can outperform SAT solvers on instances of 3-SAT problems. The NeuroSAT architecture has also been applied on special classes of crypto-analysis problems (Sun et al., 2020). In addition to supervised learning, unsupervised methods have been proposed for solving SAT problems. For Circuit-SAT a deep-gated DAG recursive neural architecture has been presented together with a differentiable training objective to optimize towards solving the Circuit-SAT problem and finding a satisfying assignment (Amizadeh et al., 2019). For Boolean satisfiability, a differentiable training objective has been proposed together with a query mechanism that allows for recurrent solution trials (Ozolins et al., 2022).
@@
In formal mathematics, deep learning has been integrated with theorem proving for premise selection (Irving et al., 2016; Wang et al., 2017; Bansal et al., 2019; Mikula et al., 2023), tactic prediction (Yang & Deng, 2019; Paliwal et al., 2020) and whole proof searches (Polu & Sutskever, 2020; First et al., 2023). For SMT formulas specifically, deep reinforcement learning has been applied to tactic prediction (Balunovic et al., 2018). In the domain of quantified boolean formulas, heuristics have been learned to guide search algorithms in proving the satisfiability and unsatisfiability of formulas (Lederman et al., 2020). For temporal logics, deep learning has been applied to prove the satisfiability of formulas and the realizability of specifications (Hahn et al., 2021; Schmitt et al., 2021; Cosler et al., 2023).
@@ 3 Proofs of (Un-)Satisfiability
We start with a brief review of certifying the (un-)satisfiability of propositional formulas in conjunctive normal form. For a set of Boolean variables $V$, we identify with each variable $x\in V$ the positive literal $x$ and the negative literal $\lnot x$ denoted by $\bar{x}$. A clause corresponds to a disjunction of literals and is abbreviated by a set of literals, e.g., $\{\bar{1},3\}$ represents ($(\lnot x_{1}\lor x_{3})$. A formula in conjunctive normal form (CNF) is a conjunction of clauses and is abbreviated by a set of clauses, e.g., $\{\{\bar{1},3\},\{1,2,\bar{4}\}\}$ represents ($(\lnot x_{1}\lor x_{3})\land(x_{1}\lor x_{2}\lor\lnot x_{4})$. Note that any Boolean formula can be converted to an equisatisfiable CNF formula in polynomial time, for example using the Tseitin transformation (Tseitin, 1983).
@@
A CNF formula is satisfiable if there exists an assignment $\mathcal{A}:V\rightarrow\{\top,\bot\}$ such that all clauses are satisfied, i.e., each clause contains a positive literal $x$ such that $\mathcal{A}(x)=\top$ or a negative literal $\bar{x}$ such that $\mathcal{A}(x)=\bot$. If no such assignment exists we call the formula unsatisfiable. To prove unsatisfiability we rely on resolution, a fundamental inference rule in satisfiability testing (Davis & Putnam, 1960). The resolution rule (Res) picks clauses with two opposite literals and performs the following inference:
@@ 4.1 General Architecture
NeuRes is a neural network that takes a CNF formula as a set of clauses and outputs either a satisfying truth assignment or a resolution proof of unsatisfiability. As such, our model comprises a formula embedder followed by an LSTM-based encoder-decoder network connected to two downstream heads: (1) an attention network responsible for selecting clause pairs, and (2) a truth assignment decoder. See Figure 1 for an overview of the NeuRes architecture. After obtaining the initial clause and literal embeddings (representing the input formula), they are passed through the encoder to get the initial hidden state of the decoder, which marks the certificate generation phase. At each step, the model selects a clause pair which gets resolved into a new clause to append to the current formula graph while decoding a candidate truth assignment in parallel. The model keeps deriving new clauses until the empty clause is found (marking resolution proof completion), a satisfying assignment is found (marking a certified sat verdict), or the limit on episode length is reached (marking timeout).
@@ 4.2 Message-Passing Embedder
Similar to NeuroSAT, we use a message-passing GNN to obtain clause and literal embeddings by performing a predetermined number of rounds. Our formula graph is also constructed in a similar fashion to NeuroSAT graphs where clause nodes are connected to their constituent literal nodes and literals are connect to their complements (cf. Appendix A). For a formula $F$ in $m$ variables and $n$ clauses, the outputs of this GNN are two matrices: $E^{L}\in\mathbb{R}^{m\times d}$ for literal embeddings and $E^{C}\in\mathbb{R}^{{n}\times d}$ for clause embedding, where $d\in\mathbb{N^{+}}$ is the embedding vector width. Here we have two key differences from NeuroSAT. Firstly, NeuroSAT uses these embeddings as voters to predict satisfiability through a classification MLP. In our case, we use these embeddings as clause tokens for clause pair selection and literal tokens for truth value assignment. Secondly, since our model derives new clauses with every resolution step, we need to embed these new clauses, as well as update existing embeddings to reflect their relation to the newly inferred clauses. Consequently, we need to introduce a new phase to the message-passing protocol, for which we explore two approaches: static embeddings and dynamic embeddings.
@@ 4.3 Selector Networks
After producing clause and literal embeddings followed by the encoder phase, NeuRes enters the derivation stage. At each step, our model needs to select two clauses to resolve, produce the resultant clause, and add it to the current formula. We tokenize clauses by their embeddings as opposed to their encoder outputs as in a Pointer Network. The reason for that is that our embeddings are dynamic, and so the encoder outputs need to be dynamic, too. While embeddings can be updated in parallel, encoder outputs are generated in sequence, which is costly.
To realize our clause-pair selection mechanism, we employ three attention-based designs. With the decoder as the main driver for our model, at each step $t$ we use the decoder hidden state vector $h_{t}$ for querying our attention networks over the set of tokens (clauses or literals).
4.3.1 Cascaded Pointer-Attention (Casc-Attn)
In this design, pairs are selected by making two consecutive attention queries on the clause pool. We condition the second attention query on the outcome (i.e., the clause) of the first query. Figure 2 shows this scheme where we perform the first query using $h_{t}$ concatenated with a zero token vector while performing the second query using $h_{t}$ concatenated with the embedding vector $E^{C}_{i}$ of the clause selected in the first query.
@@ 4.3.1 Cascaded Pointer-Attention (Casc-Attn)
Formally, Casc-Attn selects a clause index pair $(c_{1},c_{2})$ as follows:
$Q_{r}=\begin{cases}h_{t}\mathbin{\|}\mathbf{0}&\text{if }r=1\\ h_{t}\mathbin{\|}E^{C}_{c_{1}}&\text{if }r=2\end{cases}$
$c_{r}=\operatorname*{\textbf{argmax}}_{i}\left[u^{T}\tanh(W_{1}Q_{r}+W_{2}E^{C}_{i})\right]$ (2)
@@ 4.3.2 Full Self-Attention (Full-Attn)
To address the downside of independent clause selection, this variant performs self-attention between all clauses to obtain a matrix $S\in\mathbb{R}^{n\times n}$ where $S_{i,j}$ represents the attention score of the clause pair $(C_{i},C_{j})$ as shown in Figure 3. The model selects clause pairs by choosing the cell with the maximal score. In this attention scheme, clause embeddings are used as keys while queries are constructed by concatenating the decoder hidden state $h_{t}$ with all clause embeddings.
@@
$Q_{r}=W_{Q}(h_{t}\mathbin{\|}E^{C}_{r});\quad K_{r}=W_{K}E^{C}_{r};\quad S=\frac{QK^{T}}{\sqrt{d}}$
$(c_{1},c_{2})=\operatorname*{\textbf{argmax}}_{(i,j)}\,\,S_{i,j}$ (3)
where $W_{Q}\in\mathbb{R}^{2d\times d},W_{K}\in\mathbb{R}^{d\times d}$ are trainable network parameters. Since $S$ contains many cells that correspond to invalid resolution steps (i.e., clause pairs that cannot be resolved), we mask out the invalid cells from the attention grid in ensure the network selection is valid at every step.
@@ 4.3.3 Anchored Self-Attention (Anch-Attn)
In Full-Attn, the attention grid grows quadratically with the number of clauses. In this variant, we relax this cost by exploiting a property of binary resolution where each step targets a single variable in the two resolvent clauses. This allows us to narrow down candidate clause pairs by first selecting a variable as an anchor on which our clauses should be resolved. As such, we do not need to consider the full clause set at once, only the clauses containing the chosen variable $v$. We further compress the attention grid by lining clauses containing the literal $v$ on rows while lining clauses containing the literal $\neg v$ on columns. This reduces the redundancy of the attention grid since clauses containing the variable $v$ with the same parity cannot be resolved on $v$, so there is no point in matching them. In the worst case, this relaxed grid is of size $\frac{n}{2}\times\frac{n}{2}=\frac{n^{2}}{4}$ instead of $n^{2}$. In this scheme, we have two attention modules: one pointer-attention network to choose an anchor variable followed by a self-attention network to produce the anchored score grid.
@@
In light of Figure 4, this approach combines structural elements from Casc-Attn (pointer attention) and Full-Attn (self-attention); however, both elements are used differently in Anch-Attn. Firstly, pointer-attention in Casc-Attn is used to select clauses whereas Anch-Attn uses it to select variables. Secondly, self-attention in Full-Attn matches any pair of clauses ($C_{i}$, $C_{j}$) in both directions as the row and column dimensions in the attention score grid reflect the same clauses (all clauses). By contrast, Anch-Attn computes self-attention scores for clause pairs in only one order (positive instance to negative instance). Formally, Anch-Attn selects an anchor variable $v$ as follows:
$v=\operatorname*{\textbf{argmax}}\limits_{i}\left[u^{T}\tanh(W_{1}h_{t}+W_{2}(E^{L^{+}}_{i}+E^{L^{-}}_{i}))\right]$

### v2

# Learning Better Representations From Less Data For Propositional Satisfiability

@@ 1 Introduction
Boolean satisfiability (SAT) is a fundamental problem in computer science. For theory, this stems from SAT being the first problem proven NP-complete [12]. For practice, this is due to many highly-optimized SAT solvers being used as flexible reasoning engines in a variety of tasks such as model checking [11, 46], software verification [16], planning [26], and mathematical proof search [23]. Recently, SAT has also served as a litmus test for assessing the symbolic reasoning capabilities of neural models and a promising domain for neuro-symbolic systems [42, 41, 1, 9, 35]. So far, neural models only provide limited, if any, justification for unsatisfiability predictions. NeuroCore [41], for example, predicts an unsatisfiable core, the verification of which can be as hard as solving the original problem. No certificates at all or certificates that are hard to check limit neural methods in a domain where correctness is critical and prevents close integrations with symbolic methods. Therefore, we propose a neuro-symbolic model that utilizes resolution to solve SAT problems by generating easy-to-check certificates.
@@
A resolution proof is a sequence of case distinctions, each involving two clauses, that ends in the empty clause (falsum). This technique can also be used to prove satisfiability by exhaustively applying it until no further new resolution steps are possible and the empty clause has not been derived. Generating such a proof is an interesting problem from a neuro-symbolic perspective because unlike other discrete combinatorial problems that have been considered before [45, 6, 27, 29, 10], it requires selecting compatible pairs of clauses from the dynamically growing pool, as newly derived clauses are naturally considered for derivation in subsequent steps. In this work, we devise three attention-based mechanisms to perform this pair-selection needed for generating resolution proofs. In addition, we augment the architecture to efficiently handle sat (satisfiable) formulas with an assignment decoding mechanism that assigns a truth value to each literal. We hypothesize that, despite their final goals being in complete opposition, resolution and sat assignment finding can form a mutually beneficial collaboration. On the one hand, clauses derived by resolution incrementally inject additional information into the network, e.g., deriving a single-literal clause by resolution directly implies that literal should be true in any possible sat assignment. On the other hand, finding a sat assignment absolves the resolution network from having to prove satisfiability by exhaustion. On that basis, given an input formula, NeuRes proceeds in two parallel tracks: (1) finding a sat assignment, and (2) deriving a resolution proof of unsatisfiability. Both tracks operate on a shared representation of the problem state. Depending on which track succeeds, NeuRes produces the corresponding SAT verdict which is guaranteed to be sound by virtue of its certificate-based design. Since both of our certificate types are efficient to check, we can afford to perform these symbolic checks at each step. When comparing NeuRes with NeuroSAT [42], which has been trained to predict satisfiability with millions of samples, we demonstrate that NeuRes achieves a higher accuracy while providing a proof and requires only thousands of training samples.
As for most problems in theorem proving we are not only interest in finding any proof but a short proof. Resolution proofs can vary largely in their size depending on the resolution steps taken. Being able to efficiently check the proof, also allows us to adapt the proof target while training the model. In particular, we explore an expert iteration mechanism [38] that pre-rolls the resolution proof of the model and replaces the target proof whenever the pre-rolled proof is shorter. We demonstrate that the bootstrapping mechanism iteratively shortens the proofs of our training dataset while further improving the overall performance of the model.
@@
    We introduce novel architectures which combine graph neural networks with attention mechanisms for generating resolution proofs and assignments for CNF formulas (Section 4).
@@
    We show that for propositional logic, learning to prove rather than predict satisfiability results in better representations and requires far less training samples (Section 6 and 7).
@@
    We devise a bootstrapped training procedure where our model progressively produces shorter resolution proofs than its teacher (Section 6.2) boosting the model’s overall performance.
@@ 2 Related Work
SAT Solving and Certificates.
We refer to the annual SAT competitions [4] for a comprehensive overview on the ever-evolving landscape of SAT solvers, benchmarks, and proof checkers. SAT solvers are complex systems with a documented history of bugs [8, 25], hence proof certificates have been partially required in this competition since 2013 [2]. Unlike satisfiable formulas, there are several ways to certify unsatisfiable formulas [22]. Resolution proofs [50, 19] are easy to verify [14], but non-trivial to generate from modern solvers based on the paradigm of conflict-driven clause learning [33]. Clausal proofs, e.g., in DRAT format [48], are easier to generate and space-efficient, but hard to validate. Verifying the proofs can take longer than their discovery [21] and requires highly optimized algorithms [30].
@@
NeuroSAT [42] was the first study of the Boolean satisfiability problem as an end-to-end learning task. Building upon the NeuroSAT architecture, a simplified version has been trained to predict unsatisfiable cores and successfully integrated in a state-of-the-art SAT solver [41]. It has been showed that both the NeuroSAT architecture and a newly introduced deep exchangeable architecture can outperform SAT solvers on instances of 3-SAT problems [9]. The NeuroSAT architecture has also been applied on special classes of crypto-analysis problems [43]. In addition to supervised learning, unsupervised methods have been proposed for solving SAT problems. For Circuit-SAT a deep-gated DAG recursive neural architecture has been presented together with a differentiable training objective to optimize towards solving the Circuit-SAT problem and finding a satisfying assignment [1]. For Boolean satisfiability, a differentiable training objective has been proposed together with a query mechanism that allows for recurrent solution trials [35].
@@
In formal mathematics, deep learning has been integrated with theorem proving for clause selection [32, 17], premise selection [24, 47, 5, 34], tactic prediction [49, 36] and whole proof searches [39, 18]. For SMT formulas specifically, deep reinforcement learning has been applied to tactic prediction [3]. In the domain of quantified boolean formulas, heuristics have been learned to guide search algorithms in proving the satisfiability and unsatisfiability of formulas [31]. For temporal logics, deep learning has been applied to prove the satisfiability of linear-time temporal logic formulas and the realizability of specifications [20, 40, 13].
@@ 3 Proofs of (Un-)Satisfiability
We start with a brief review of certifying the (un-)satisfiability of propositional formulas in conjunctive normal form. For a set of Boolean variables $V$, we identify with each variable $x\in V$ the positive literal $x$ and the negative literal $\lnot x$ denoted by $\bar{x}$. A clause corresponds to a disjunction of literals and is abbreviated by a set of literals, e.g., $\{\bar{1},3\}$ represents ($(\lnot x_{1}\lor x_{3})$. A formula in conjunctive normal form (CNF) is a conjunction of clauses and is abbreviated by a set of clauses, e.g., $\{\{\bar{1},3\},\{1,2,\bar{4}\}\}$ represents ($(\lnot x_{1}\lor x_{3})\land(x_{1}\lor x_{2}\lor\lnot x_{4})$. Any Boolean formula can be converted to an equisatisfiable CNF formula in polynomial time, for example with Tseitin transformation [44].
@@
A CNF formula is satisfiable if there exists an assignment $\mathcal{A}:V\rightarrow\{\top,\bot\}$ such that all clauses are satisfied, i.e., each clause contains a positive literal $x$ such that $\mathcal{A}(x)=\top$ or a negative literal $\bar{x}$ such that $\mathcal{A}(x)=\bot$. If no such assignment exists we call the formula unsatisfiable. To prove unsatisfiability we rely on resolution, a fundamental inference rule in satisfiability testing [15]. The resolution rule (Res) picks clauses with two opposite literals and performs the following inference:
@@ 4.1 General Architecture
NeuRes is a neural network that takes a CNF formula as a set of clauses and outputs either a satisfying truth assignment or a resolution proof of unsatisfiability. As such, our model comprises a formula embedder connected to two downstream heads: (1) an attention network responsible for selecting clause pairs, and (2) a truth assignment decoder. See Figure 1 for an overview of the NeuRes architecture. After obtaining the initial clause and literal embeddings (representing the input formula), we continue with the iterative certificate generation phase. At each step, the model selects a clause pair which gets resolved into a new clause to append to the current formula graph while decoding a candidate truth assignment in parallel. The model keeps deriving new clauses until the empty clause is found (marking resolution proof completion), a satisfying assignment is found (marking a certified sat verdict), or the limit on episode length is reached (marking timeout).
@@ 4.2 Message-Passing Embedder
Similar to NeuroSAT, we use a message-passing GNN to obtain clause and literal embeddings by performing a predetermined number of rounds. Our formula graph is also constructed in a similar fashion to NeuroSAT graphs where clause nodes are connected to their constituent literal nodes and literals are connect to their complements (cf. Appendix A). For a formula $F$ in $m$ variables and $n$ clauses, the outputs of this GNN are two matrices: $E^{L}\in\mathbb{R}^{m\times d}$ for literal embeddings and $E^{C}\in\mathbb{R}^{{n}\times d}$ for clause embedding, where $d\in\mathbb{N^{+}}$ is the embedding dimension. Here we have two key differences from NeuroSAT. Firstly, NeuroSAT uses these embeddings as voters to predict satisfiability through a classification MLP. In our case, we use these embeddings as clause tokens for clause pair selection and literal tokens for truth value assignment. Secondly, since our model derives new clauses with every resolution step, we need to embed these new clauses, as well as update existing embeddings to reflect their relation to the newly inferred clauses. Consequently, we need to introduce a new phase to the message-passing protocol, for which we explore two approaches: static embeddings and dynamic embeddings.
@@ 4.3 Selector Networks
After producing clause and literal embeddings, NeuRes enters the derivation stage. At each step, our model needs to select two clauses to resolve, produce the resultant clause, and add it to the current formula. To realize our clause-pair selection mechanism, we employ three attention-based designs.
4.3.1 Cascaded Attention (Casc-Attn)
@@
In this design, pairs are selected by making two consecutive attention queries on the clause pool. We condition the second attention query on the outcome (i.e., the clause) of the first query. Figure 2 shows this scheme where we perform the first query using the mean of the literal embeddings $\overline{E^{L}}$ concatenated with a zero vector while performing the second query using the mean of the literal embeddings concatenated with the embedding vector $E^{C}_{c_{1}}$ of the clause selected in the first query. Formally, Casc-Attn selects a clause index pair $(c_{1},c_{2})$ as follows:
$c_{i}=\operatorname*{\textbf{argmax}}_{j}\left[u^{T}\tanh(W_{1}q_{i}+W_{2}E^{C}_{j})\right]\quad\text{with}\quad q_{i}=\begin{cases}\overline{E^{L}}\mathbin{\|}\mathbf{0}&\text{if }i=1\\ \overline{E^{L}}\mathbin{\|}E^{C}_{c_{1}}&\text{if }i=2\end{cases}$ (1)
@@ 4.3.2 Full Self-Attention (Full-Attn)
To address the downside of independent clause selection, this variant performs self-attention between all clauses to obtain a matrix $S\in\mathbb{R}^{n\times n}$ where $S_{i,j}$ represents the attention score of the clause pair $(c_{i},c_{j})$ as shown in Figure 3. The model selects clause pairs by choosing the cell with the maximal score. In this attention scheme, the clause embeddings are used as both queries and keys.
@@
$(c_{1},c_{2})=\operatorname*{\textbf{argmax}}_{(i,j)}\,\,S_{i,j}\quad\text{with}\quad Q=E^{C}W_{Q};\quad K=E^{C}W_{K};\quad S=\frac{QK^{T}}{\sqrt{d}}$ (2)
where $W_{Q}\in\mathbb{R}^{d\times d},W_{K}\in\mathbb{R}^{d\times d}$ are trainable network parameters. Since $S$ contains many cells that correspond to invalid resolution steps (i.e., clause pairs that cannot be resolved), we mask out the invalid cells from the attention grid in ensure the network selection is valid at every step.
@@ 4.3.3 Anchored Self-Attention (Anch-Attn)
In Full-Attn, the attention grid grows quadratically with the number of clauses. In this variant, we relax this cost by exploiting a property of binary resolution where each step targets a single variable in the two resolvent clauses. This allows us to narrow down candidate clause pairs by first selecting a variable as an anchor on which our clauses should be resolved. As such, we do not need to consider the full clause set at once, only the clauses containing the chosen variable $v$. We further compress the attention grid by lining clauses containing the literal $v$ on rows while lining clauses containing the literal $\neg v$ on columns. This reduces the redundancy of the attention grid since clauses containing the variable $v$ with the same parity cannot be resolved on $v$, so there is no point in matching them. In the worst case, this relaxed grid is of size $\frac{n}{2}\times\frac{n}{2}=\frac{n^{2}}{4}$ instead of $n^{2}$. In this scheme, we have two attention modules: one attention network to choose an anchor variable followed by a self-attention network to produce the anchored score grid.
@@
In light of Figure 4, this approach combines structural elements from Casc-Attn and Full-Attn; however, both elements are used differently in Anch-Attn. Firstly, the attention mechanism in Casc-Attn is used to select clauses whereas Anch-Attn uses it to select variables. Secondly, self-attention in Full-Attn matches any pair of clauses ($c_{i}$, $c_{j}$) in both directions as the row and column dimensions in the attention score grid reflect the same clauses (all clauses). By contrast, Anch-Attn computes self-attention scores for clause pairs in only one order (positive instance to negative instance). Formally, Anch-Attn selects an anchor variable $v$ as follows:
$v=\operatorname*{\textbf{argmax}}\limits_{i}\left[u^{T}\tanh(W_{1}\overline{E^{C}}+W_{2}(E^{L^{+}}_{i}+E^{L^{-}}_{i}))\right]$ (3)